# LAMP: Data-Efficient Linear Affine Weight-Space Models for Parameter-Controlled 3D Shape Generation and Extrapolation

### v1

@@ 1 Introduction
Many existing methods achieve control by learning a shared latent space and then traversing or disentangling latent dimensions to align with semantic or parametric factors (Vahdat et al., 2022; Zhao et al., 2023; Morita et al., 2024; Xiang et al., 2025). While powerful, latent approaches require a single generator to capture all variability, achieving clean, interpretable directions often proves difficult. The problem is particularly challenging in 3D shape generation, as most 3D generative models—whether voxel, point cloud, mesh, or implicit—rely on large datasets and provide few guarantees about parameter control or generalization beyond the training distribution.
@@ 2 Related Work
A wide range of 3D generative models have been proposed, spanning voxel grids Wu & et al. (2016), point clouds Achlioptas & et al. (2018), meshes Groueix & et al. (2018), and neural implicit representations such as signed distance functions (SDFs) Park et al. (2019); Chibane et al. (2020). Neural SDFs capture high-resolution geometry and have been applied to generation and reconstruction Chen & et al. (2022) and reconstruction Atzmon & Lipman (2020). Recent diffusion-based 3D generators operate on implicit or latent representations, including LION, GET3D, Diffusion-SDF, SDFusion, and SALAD Zeng et al. (2022); Gao et al. (2022); Chou et al. (2023); Cheng et al. (2023); Koo et al. (2023). However, these methods typically assume abundant training data and lack explicit mechanisms for parameter-constrained generation or safe extrapolation. More recently, HyperDiffusion Erkoç et al. (2023) modeled the distribution of overfit implicit networks directly in weight space, sampling new fields via diffusion. These works treat networks themselves as data points in parameter space, but focus on unconditional sampling or learned meta-combination rather than explicit affine mixing under interpretable constraints.
@@
Efforts to introduce control often rely on conditioning on labels or attributes Gao et al. (2019); Niemeyer et al. (2020), or on parametric templates derived from CAD data Yumer & Mitra (2016); Wang et al. (2022). Diffusion models have recently been adapted for class-conditional and partially conditional shape generation Chen & et al. (2023); Liu & et al. (2023). Multimodal conditioning has also been explored for controllable 3D generation, e.g., CLIP-Forge and Michelangelo Sanghi et al. (2022); Zhao et al. (2023). However, these methods rarely support precise parameter specification (e.g., generating a car body with fixed ramp angle and width) and are not designed for data-efficient regimes. Our approach complements this line of work by enabling direct control through interpretable parameters.
@@
Engineering design relies heavily on parametric modeling, where small sets of interpretable variables govern global shape Seff & et al. (2020); Wang et al. (2022); Du & et al. (2023). While recent learning-based CAD systems leverage symbolic histories or constraint graphs, they often require large, structured datasets. By contrast, we target the data-efficient setting where only meshes and parameter annotations are available. LAMP directly links parameters to mesh geometry via aligned SDF weight spaces, and enforces validity through a linearity-mismatch safety metric and mesh-based surrogate checks.
@@
Latent-space interpolation has been widely explored in autoencoders Achlioptas & et al. (2018) and implicit representations Park et al. (2019), but these spaces are often not semantically aligned, leading to unrealistic interpolations or invalid extrapolations. Our method leverages affine mixing of aligned SDF weights, which—combined with the linearity-mismatch criterion—ensures that interpolated or extrapolated meshes remain geometrically consistent and satisfy parametric constraints.
@@
Our approach bridges these threads. Like HyperDiffusion, we treat overfit exemplar networks as aligned points in parameter space, but instead of learning a generative model over them, we provide direct, interpretable control by solving for mixing coefficients in parameter space and applying them in weight space. Unlike latent traversal or disentanglement, we do not rely on a single model to encode all variation. And unlike prior weight interpolation methods, we generalize beyond two-model blends to a bank of exemplars, enabling constraint-driven, multi-way affine mixing. To our knowledge, this is the first work to formulate controllable generation as parameter-space affine mixing of aligned exemplar networks.
@@ 3 Method
An arbitrary weight vector $\mathbf{w}$ is decoded into a mesh $\mathcal{M}=\mathrm{Decode}(\mathbf{w})$ by evaluating the zero-level set of the SDF distribution on a dense voxel grid at the desired resolution, and extracting the isosurface using the marching cubes algorithm Lorensen & Cline (1998).
@@ 4 Experiments, Results, and Discussion
We benchmark these methods on two recent parametric datasets (Appendix I): DrivAerNet++: a large-scale multimodal car dataset with $\sim 8{,}000$ distinct geometries, each annotated with $26$ interpretable design parameters, high-resolution meshes, and CFD-based aerodynamic coefficients. Elrefaie et al. (2024) BlendedNet: a blended wing-body (BWB) aircraft dataset with 999 geometries, each simulated under $9$ flight conditions, and annotated with planform parameters such as chord-length ratios, spanwise widths, and sweep angles. Sung et al. (2025) For brevity, we focus on DrivAerNet++ in the main paper and report full BlendedNet results in Appendix D.
@@
For in-dataset generation (when the target mesh is available), we evaluate: Chamfer Distance (CD) and Intersection-over-Union (IoU) between generated and reference meshes (see Appendix F). Parameter Error: mean absolute error (MAE) between target parameters and surrogate-predicted parameters for the generated mesh. For out-of-distribution extrapolation (no ground-truth mesh), we evaluate: Parameter Fidelity:surrogate-predicted MAE and $R^{2}$ between target parameters $p_{d}$ and inferred $\hat{p}$, Distributional Similarity:Minimum Matching Distance (MMD) between generated shapes and a reference set of geometries. (see Appendix F).
@@
To assess whether generated shapes respect design-parameter constraints, we employ a mesh-based surrogate model trained to predict interpretable parameters (e.g., geometric or performance attributes). We compute fixed, randomly initialized PointNet embeddings for each decoded mesh Amid et al. (2022), and fit a LASSO regressor Tibshirani (1996) to map embeddings to physical parameters. Despite using an untrained encoder, the surrogate consistently achieves $R^{2}>0.9$ on held-out test sets, providing a robust parameter validator. When possible, we further cross-check compliance through direct geometric measurements of the meshes (details in Appendix G).
@@ Results.
We validate this metric against a human-annotated dataset of valid and invalid meshes (Appendix H). The results show excellent discriminative power (ROC AUC = 0.989, PR AUC = 0.990), and $\epsilon=0.01$ emerges as a reliable threshold for separating valid from invalid generations. Additional examples of failure cases and validation analysis are provided in Figs. 1718.
Another limitation is that we were only able to hard-code the measurement of car length, as explained in Appendix G, achieving $R^{2}=0.999$ under $\pm 100\%$ extrapolation sweeps. In contrast, other parameters were far more difficult to hard-code due to the complexity of their deformations across different car geometries and the absence of reliable methods for accurate measurement.
@@
Figure 5: Linearity-mismatch safety metric for diffuser angle extrapolation. Failures (e.g., sample f) occur when the metric exceeds the threshold. See Appendix H for more.
@@
To validate the results, we use the mesh-based surrogate (Appendix  G) to predict both physical parameters and drag coefficients from the generated meshes. The surrogate predictions are compared to ground-truth values, with predicted vs. target plots provided in the Appendix. We evaluate two objectives: (i) parameter fidelity, i.e. how closely the generated meshes respect the selected physical parameter constraints, and (ii) drag fidelity, i.e. how accurately the achieved reduction matches the $10\%$ target. Here, decay MAE denotes the mean absolute error between the desired $10\%$ decay and the observed (predicted) decay, averaged across all samples.
@@ 5 Conclusion
We presented LAMP, a data-efficient framework for parameter-controlled 3D mesh generation that leverages affine mixing in aligned SDF weight spaces and a linearity-based safety metric. Experiments on DrivAerNet++ and BlendedNet show that LAMP outperforms conditional autoencoders and DNI across interpolation, large-range extrapolation, and performance-guided optimization, achieving reliable control with as few as 100 exemplars. The safety score provides a principled safeguard in low-data regimes, addressing a key challenge for robust generalization. LAMP advances the goal of controllable, efficient, and verifiable 3D generation, common in engineering applications. Future directions include extending the framework to partial parameter observability, multi-modal conditioning, and integrating physics-aware constraints.
@@ Appendix E Ablation Study: How Does Sample Size Affect Reliability and Extrapolation in LAMP?
Appendix F Evaluation Metrics
@@
Appendix G Constraint Compliance Validation: Surrogates and Direct Measurements
@@
Figure 12 illustrates the surrogate pipeline. Figures 14 and 13 show predicted versus ground-truth parameter values on the BlendedNet and DrivAerNet++ datasets, respectively, demonstrating high accuracy across test sets. We train the surrogates on $800$ samples and evaluate on a held-out test set of $200$ samples.
@@
Appendix H Validation of the Linearity-Mismatch Metric Against Human Annotated Data
@@ Discussion.
These experiments demonstrate that the linearity-mismatch metric is a reliable quantitative proxy for mesh validity. Its strong agreement with human annotations justifies our use of $\epsilon=0.01$ as the default safety threshold throughout this work.
@@
Appendix I Parametrization of DrivAerNet++ and BlendedNet
@@
Appendix J LLM Usage

### v2

@@ 1 Introduction
Many existing methods achieve control by learning a shared latent space and then traversing or disentangling latent dimensions to align with semantic or parametric factors (Vahdat et al., 2022; Zhao et al., 2023; Morita et al., 2024; Xiang et al., 2025). While powerful, latent approaches require a single generator to capture all variability, achieving clean, interpretable directions often proves difficult. The problem is particularly challenging in 3D shape generation, as most 3D generative models, whether voxel, point cloud, mesh, or implicit, rely on large datasets and provide few guarantees about parameter control or generalization beyond the training distribution.
@@ 2 Related Work
A wide range of 3D generative models have been proposed, spanning voxel grids Wu et al. (2016), point clouds Achlioptas et al. (2018), meshes Groueix et al. (2018), and neural implicit representations such as signed distance functions (SDFs) Park et al. (2019); Chibane et al. (2020). Neural SDFs capture high-resolution geometry and have been applied to generation and reconstruction Jiang et al. (2020) and reconstruction Atzmon and Lipman (2020). Recent diffusion-based 3D generators operate on implicit or latent representations, including LION, GET3D, Diffusion-SDF, SDFusion, and SALAD Vahdat et al. (2022); Gao et al. (2022); Chou et al. (2023); Cheng et al. (2023); Koo et al. (2023). However, these methods typically assume abundant training data and lack explicit mechanisms for parameter-constrained generation or safe extrapolation. More recently, HyperDiffusion Erkoç et al. (2023) modeled the distribution of overfit implicit networks directly in weight space, sampling new fields via diffusion. These works treat networks themselves as data points in parameter space, but focus on unconditional sampling or learned meta-combination rather than explicit affine mixing under interpretable constraints.
@@
Efforts to introduce control often rely on conditioning on labels or attributes Gao et al. (2019); Niemeyer et al. (2020), or on parametric templates derived from CAD data Yumer and Mitra (2016); Wu et al. (2021). Diffusion models have recently been adapted for class-conditional and partially conditional shape generation Vahdat et al. (2022). Multimodal conditioning has also been explored for controllable 3D generation, e.g., CLIP-Forge and Michelangelo Sanghi et al. (2022); Zhao et al. (2023). However, these methods rarely support precise parameter specification (e.g., generating a car body with fixed ramp angle and width) and are not designed for data-efficient regimes. Our approach complements this line of work by enabling direct control through interpretable parameters.
@@
Engineering design relies heavily on parametric modeling, where small sets of interpretable variables govern global shape Seff et al. (2020); Wu et al. (2021). While recent learning-based CAD systems leverage symbolic histories or constraint graphs, they often require large, structured datasets. By contrast, we target the data-efficient setting where only meshes and parameter annotations are available. LAMP directly links parameters to mesh geometry via aligned SDF weight spaces, and enforces validity through a linearity-mismatch safety metric and mesh-based surrogate checks.
@@
Latent-space interpolation has been widely explored in autoencoders Achlioptas et al. (2018) and implicit representations Park et al. (2019), but these spaces are often not semantically aligned, leading to unrealistic interpolations or invalid extrapolations. Our method leverages affine mixing of aligned SDF weights, which, combined with the linearity-mismatch criterion, ensures that interpolated or extrapolated meshes remain geometrically consistent and satisfy parametric constraints.
@@
Our approach bridges these threads. Like HyperDiffusion Erkoç et al. (2023), we treat overfit exemplar networks as aligned points in parameter space, but instead of learning a generative model over them, we provide direct, interpretable control by solving for mixing coefficients in parameter space and applying them in weight space. Unlike latent traversal or disentanglement, we do not rely on a single model to encode all variation. And unlike prior weight interpolation methods, we generalize beyond two-model blends to a bank of exemplars, enabling constraint-driven, multi-way affine mixing. To our knowledge, this is the first work to formulate controllable generation as parameter-space affine mixing of aligned exemplar networks.
@@ 3 Method
An arbitrary weight vector $\mathbf{w}$ is decoded into a mesh $\mathcal{M}=\mathrm{Decode}(\mathbf{w})$ by evaluating the zero-level set of the SDF distribution on a dense voxel grid at the desired resolution, and extracting the isosurface using the marching cubes algorithm Lorensen and Cline (1998).
@@ 4 Experiments, Results, and Discussion
We benchmark these methods on two recent parametric datasets (Appendix J): DrivAerNet++: a large-scale multimodal car dataset with $\sim 8{,}000$ distinct geometries, each annotated with $26$ interpretable design parameters, high-resolution meshes, and CFD-based aerodynamic coefficients. Elrefaie et al. (2024) BlendedNet: a blended wing-body (BWB) aircraft dataset with 999 geometries, each simulated under $9$ flight conditions, and annotated with planform parameters such as chord-length ratios, spanwise widths, and sweep angles. Sung et al. (2025) For brevity, we focus on DrivAerNet++ in the main paper and report full BlendedNet results in Appendix D.
@@
For in-dataset generation (when the target mesh is available), we evaluate: Chamfer Distance (CD) and Intersection-over-Union (IoU) between generated and reference meshes (see Appendix G). Parameter Error: mean absolute error (MAE) between target parameters and surrogate-predicted parameters for the generated mesh. For out-of-distribution extrapolation (no ground-truth mesh), we evaluate: Parameter Fidelity:surrogate-predicted MAE and $R^{2}$ between target parameters $p_{d}$ and inferred $\hat{p}$, Distributional Similarity:Minimum Matching Distance (MMD) between generated shapes and a reference set of geometries. (see Appendix G).
@@
To assess whether generated shapes respect design-parameter constraints, we employ a mesh-based surrogate model trained to predict interpretable parameters (e.g., geometric or performance attributes). We compute fixed, randomly initialized PointNet embeddings for each decoded mesh Amid et al. (2022), and fit a LASSO regressor Tibshirani (1996) to map embeddings to physical parameters. Despite using an untrained encoder, the surrogate consistently achieves $R^{2}>0.9$ on held-out test sets, providing a robust parameter validator. When possible, we further cross-check compliance through direct geometric measurements of the meshes (details in Appendix H).
@@ Results.
We validate this metric against a human-annotated dataset of valid and invalid meshes (Appendix I). The results show excellent discriminative power (ROC AUC = 0.989, PR AUC = 0.990), and $\epsilon=0.01$ emerges as a reliable threshold for separating valid from invalid generations. Additional examples of failure cases and validation analysis are provided in Figs. 1819.
Another limitation is that we were only able to hard-code the measurement of car length, as explained in Appendix H, achieving $R^{2}=0.999$ under $\pm 100\%$ extrapolation sweeps. In contrast, other parameters were far more difficult to hard-code due to the complexity of their deformations across different car geometries and the absence of reliable methods for accurate measurement.
@@
Figure 5: Linearity-mismatch safety metric for diffuser angle extrapolation. Failures (e.g., sample f) occur when the metric exceeds the threshold. See Appendix I for more.
@@
To validate the results, we use the mesh-based surrogate (Appendix  H) to predict both physical parameters and drag coefficients from the generated meshes. The surrogate predictions are compared to ground-truth values, with predicted vs. target plots provided in the Appendix. We evaluate two objectives: (i) parameter fidelity, i.e. how closely the generated meshes respect the selected physical parameter constraints, and (ii) drag fidelity, i.e. how accurately the achieved reduction matches the $10\%$ target. Here, decay MAE denotes the mean absolute error between the desired $10\%$ decay and the observed (predicted) decay, averaged across all samples.
@@
5 Limitations and Future Work
While LAMP enables data-efficient and interpretable parameter-controlled 3D generation, it has several limitations. First, the method assumes that all exemplars share a common topological structure and satisfy the linear control-point model in Assumption A1; using geometries with differing topology or strongly nonlinear deformations would violate this assumption and invalidate affine mixing. Second, reliable extrapolation depends on remaining within the locally linear weight-space regime, which may break under large extrapolation or insufficient exemplar coverage. Finally, LAMP may fail when target performance objectives and physical parameters impose conflicting requirements, yielding combinations that are not jointly realizable within the exemplar set. Future work will aim to extend this framework to support multiple topological families, integrate physics-aware constraints, and develop multi-modal conditioning pipelines.
@@ 6 Conclusion
We presented LAMP, a data-efficient framework for parameter-controlled 3D mesh generation that leverages affine mixing in aligned SDF weight spaces and a linearity-based safety metric. Experiments on DrivAerNet++ and BlendedNet show that LAMP outperforms conditional autoencoders and DNI across interpolation, large-range extrapolation, and performance-guided optimization, achieving reliable control with as few as 100 exemplars. The safety score provides a principled safeguard in low-data regimes, addressing a key challenge for robust generalization. LAMP advances the goal of controllable, efficient, and verifiable 3D generation, common in engineering applications.
@@ Appendix E Ablation Study: How Does Sample Size Affect Reliability and Extrapolation in LAMP?
Appendix F Ablation Study: How Does SDF Decoder Finetuning Affect Reliability and Extrapolation in LAMP?
Sensitivity to Initialization
To evaluate whether LAMP depends on a shared initialization during SDF overfitting, we trained exemplar SDF decoders both with and without a common initialization. With a shared initialization, all exemplars remain in a common local basin, preserving strong weight-space alignment and enabling smooth, parameter-consistent interpolations.
Removing the shared initialization forces each decoder to converge to a different basin, breaking weight-space alignment. Under this setting, affine interpolation produces severely distorted and non–car-like shapes, confirming that a shared initialization is essential for maintaining a coherent basis.
Robustness to Random Seeds
We next tested whether varying the random seed (while keeping the shared initialization fixed) affects weight-space alignment. Using different seeds slightly reduces performance but does not break alignment: generated shapes remain visually consistent, parameter predictions stay stable, and the linearity assumption (A2) continues to hold.
In contrast, removing the shared initialization destroys linearity entirely and leads to invalid, severely distorted generations.
To quantify reliability, we measured $R^{2}$, MAE, and safe extrapolation range over single-parameter sweeps using 100 exemplars. Results are summarized in Table 8.
Table 8: Effect of initialization and random seeds on reliability and extrapolation behavior in LAMP, evaluated over 100 exemplar SDF decoders.
Condition $R^{2}$ MAE Mean Safe Extrapolation Range
Same Seed 0.838 0.507 $330\%$
Random Seed 0.773 0.610 $304\%$
No Initialization 37.77-37.77 9.16 $0\%$
Figure 12: Representative comparison of weight-space interpolation under different training conditions: Left: shared initialization (stable, coherent interpolation). Middle: random seeds (still stable). Right: no shared initialization (interpolation collapses).
Appendix G Evaluation Metrics
@@
Appendix H Constraint Compliance Validation: Surrogates and Direct Measurements
@@
Figure 13 illustrates the surrogate pipeline. Figures 15 and 14 show predicted versus ground-truth parameter values on the BlendedNet and DrivAerNet++ datasets, respectively, demonstrating high accuracy across test sets. We train the surrogates on $800$ samples and evaluate on a held-out test set of $200$ samples.
@@
Appendix I Validation of the Linearity-Mismatch Metric Against Human Annotated Data
@@ Discussion.
These experiments demonstrate that the linearity-mismatch metric is a reliable quantitative proxy for mesh validity. Its strong agreement with human annotations justifies our use of $\epsilon=0.01$ as the default safety threshold throughout this work. Moreover, the threshold can be strict or relaxed depending on the application’s tolerance for geometric distortion. In aerodynamic design studies or performance-critical evaluations, we adopt a conservative cutoff to avoid any degradation in shape fidelity. In contrast, exploratory design settings may permit larger mismatch values to encourage diversity in the generated geometries. Practitioners may therefore choose a threshold smaller or larger than $0.01$ depending on the demands of their downstream task.
@@
Appendix J Parametrization of DrivAerNet++ and BlendedNet
@@
Appendix K Quantitative Validation of Linearity Assumption (A2)
To assess the linearity assumption (A2), we compared the signed distance field (SDF) predicted by the interpolated network,
$f\!\left(z;\sum_{i}\alpha_{i}w_{i}\right),$
against the linear combination of individual model outputs,
$\sum_{i}\alpha_{i}\,f(z;w_{i}).$
We sampled 10,000 random 3D query points and evaluated both quantities across a range of interpolation and mild extrapolation factors. As shown in Fig. 22, the relationship is nearly perfectly linear in-distribution, with $R^{2}\approx 0.99$. Linearity degrades gradually as the extrapolation factor increases.
This empirical analysis demonstrates that SDF networks exhibit approximate linear behavior under weight-space interpolation, supporting assumption (A2).
Figure 22: Linearity comparison of $f(z;\sum_{i}\alpha_{i}w_{i})$ vs. $\sum_{i}\alpha_{i}f(z;w_{i})$ evaluated over 10,000 3D query points while varying the extrapolation factor.
Appendix L Extending LAMP to Non-Implicit Decoders
To examine whether LAMP can be applied beyond implicit SDF models, we performed an experiment using a non-implicit point cloud decoder that operates by deforming a spherical template into the target point cloud. First, we overfit this deformation-based decoder to a single target shape (left). We then fine-tuned the same model on a second target shape (middle), producing two distinct sets of network weights corresponding to two different geometries.
With these two trained decoders, we linearly interpolated between their weight vectors and decoded the resulting intermediate geometry. As shown in Fig. 23, the interpolated weights produce a smooth and coherent shape that lies between the two endpoints (right), confirming that weight-space interpolation remains valid even in non-implicit architectures.
This experiment demonstrates that the LAMP principle, leveraging linearity and smoothness under weight interpolation for controllable geometry manipulation, extends naturally to deformation-based point cloud decoders. This suggests that LAMP provides a general mechanism for parameterized geometry control across diverse neural shape representations.
Figure 23: Interpolation using a deformation-based point cloud decoder. Left: model overfit to the first target (sphere deformed into shape A). Middle: model fine-tuned on the second target (sphere deformed into shape B). Right: linear interpolation in weight space yields a valid intermediate geometry, showing that LAMP extends beyond implicit SDF decoders.
Appendix M Multi-Objective LAMP: Balancing Parameter Satisfaction and Safe Generation
Formulation
A natural extension of LAMP is to treat parameter matching and geometric reliability as a multi-objective optimization. Instead of minimizing only the parameter error, we solve
$\min_{\alpha}\ \bigl\|\mathbf{P}_{:\!,\mathcal{C}}^{\top}\alpha-\mathbf{p}_{d,\mathcal{C}}\bigr\|_{2}^{2}\;+\;\lambda\,\|\alpha\|_{2}^{2}\quad\text{s.t.}\quad\mathbf{1}^{\top}\alpha=1.$ (3)
where $\|\alpha\|_{2}$ serves as a fast extrapolation proxy. Evaluating the full linearity-mismatch safety metric at each iteration requires evaluating the $N$ unflattened SDF decoders for each of the 3D query points, which is computationally heavy. In contrast, $\|\alpha\|$ varies monotonically with extrapolation and correlates strongly with the final safety score, enabling efficient exploration of accuracy vs reliability trade-offs. Since LAMP’s optimization is extremely fast, sweeping $\lambda$ yields the Pareto frontier essentially for free, letting users select conservative, balanced, or exploratory solutions depending on application tolerance.
Conflicting Parameters and Safety Behavior
Conflicting or mutually incompatible parameter targets often force the optimization outside the valid linear regime, producing large $\|\alpha\|$ and correspondingly high safety scores. The leftmost example in Fig. 24 shows such a case: although the parameter error (MAE) is low, the required extrapolation produces visible geometric distortion. When decoded into meshes, these settings exhibit collapse or deformation, and the linearity-mismatch metric correctly flags them as invalid. This multi-objective formulation provides a principled way to avoid such failure modes while still permitting controlled extrapolation for creative or exploratory design tasks.
Figure 24: Trade-off between parameter error (MAE), extrapolation ($\|\alpha\|$), and geometric safety. The leftmost point illustrates conflicting parameter constraints: despite a low MAE, the high $\|\alpha\|$ and safety metric indicate that the solution lies outside the valid linear region, leading to distorted geometries after decoding.
Appendix N Quantifying LAMP’s High-Dimensional Extrapolation Volume
To quantify how LAMP expands the reachable design space under controlled extrapolation, we perform the analysis on a subset of 23 parameters. Each selected parameter is scaled to lie in [1,1][-1,1], and extrapolation corresponds to enlarging this 23-dimensional hypercube to
$1-f,\;1+f]^{23},$
which has volume $(1+f)^{23}$ relative to the original domain.
Monte Carlo Approximation
For each extrapolation factor $f$, we uniformly sample 10,00010{,}000 random parameter vectors in the expanded domain $1-f,\,1+f]^{23}$. For each sample, we solve for mixture coefficients using LAMP and generate the corresponding mesh. We then evaluate the linearity-mismatch safety metric to determine whether the geometry is valid.
Let $\text{Valid}(f)$ denote the fraction of samples whose meshes pass the safety threshold. The extrapolation volume ratio for factor $f$ is then estimated as
$\text{VolRatio}(f)\;=\;(1+f)^{23}\times\text{Valid}(f),$
providing an aggregate measure of how much of the enlarged parameter space remains safely reachable.
Results
LAMP provides reliable parameter control even under extrapolation. Using only 100 normalized exemplar shapes, LAMP already achieves high-accuracy mixing and maintains strong local linearity. As shown in Fig. 25(b), extrapolating each parameter by just 50% expands the reachable design space by a factor of $\sim 8000\times$ while still preserving a high validity ratio under the safety metric. This demonstrates that modest extrapolation combined with sufficient coverage of the parameter space enables both accurate control and large-scale geometry exploration.
(a) Validity ratio vs. extrapolation factor.
(b) Extrapolation volume ratio.
Figure 25: (a) Validity ratio under increasing extrapolation factor $f$. (b) Corresponding extrapolation-volume ratio in the 23-D normalized parameter space. LAMP achieves accurate control with only 100 samples and, with a 50% extrapolation ($f=0.5$), enables exploration of a design space approximately 8000× larger than the dataset volume.
Appendix O Mixing Behavior on Geometries with Topological Features
To assess whether LAMP can reliably handle shapes that include topological features such as holes, we conducted a controlled study on a simple but representative part: a rectangular plate with a circular through-hole. The geometry is parameterized by six parameters: plate length, width, and thickness, and hole radius, and hole x- and y-positions.
We trained a set of $13$ exemplar SDF decoders from a shared initialization and applied LAMP to generate new shapes by varying each parameter independently. For every parameter, we performed a sweep that extends beyond the training range by $\pm 100\%$ of the original dataset bounds.
As shown in Fig. 26, LAMP successfully produces valid, smoothly varying geometries throughout these extended sweeps. In particular, the hole feature remains stable under extrapolation, and its radius and position vary as expected even when moved well outside the dataset range. These results demonstrate that LAMP is capable of mixing weights for shapes that contain localized geometric and topological features, maintaining structural coherence even under aggressive extrapolation.
Figure 26: Parameter sweeps on a plate-with-hole geometry using LAMP. Each column (A–E) corresponds to normalized parameter values extending to $\pm 100\%$ beyond the dataset range. LAMP maintains coherent geometry and stable topological structure across all extrapolations.
Appendix P LLM Usage